Solve: m-5=1/2(12-14m)
Answer by jim_thompson5910(35256)

m-5=1/2(12-14m)

2(m-5)=12-14m

2m-10=12-14m

2m=12-14m+10

2m+14m=12+10

16m=12+10

16m=22

m=(22)/(16)

m=11/8

Answer: m = 11/8